Complete each statement and Round to the nearest hundredth.
72 in. = _________ cm
Help not good with this kind of stuff at all!!

Answer by checkley71(8403)

1 in.=2.54 cm.
1/2.54=72/x now cross multiply.
x=2.54*72
x=182.88 cm is the answer for converting 72 inches to cm.